The probability that a year selected at random will have $53$ Mondays is

  • A
    $\frac{1}{4}$
  • B
    $\frac{3}{28}$
  • C
    $\frac{5}{28}$
  • D
    $\frac{3}{4}$

Explore More

Similar Questions

The probability that the product of the outcomes when three dice are rolled simultaneously is divisible by $4$ is equal to

Of the three independent events $E_1, E_2$ and $E_3$,the probability that only $E_1$ occurs is $\alpha$,only $E_2$ occurs is $\beta$ and only $E_3$ occurs is $\gamma$. Let the probability $p$ that none of events $E_1, E_2$ or $E_3$ occurs satisfy the equations $(\alpha - 2\beta)p = \alpha\beta$ and $(\beta - 3\gamma)p = 2\beta\gamma$. All the given probabilities are assumed to lie in the interval $(0, 1)$. Then $\frac{\text{Probability of occurrence of } E_1}{\text{Probability of occurrence of } E_3} = $

If $S$ is the sample space of a random experiment $\xi$ and $P$ is a probability function defined on the power set $\mathcal{P}(S)$ of $S$,then which one of the following is not satisfied by $P$?
$(i)$ $P(\phi) = 0$
(ii) If $E^c$ is the complementary event of $E$,then $P(E^c) = 1 - P(E)$
(iii) $0 \leq P(E) \leq 1, \forall E \subseteq S$
(iv) If $E_1 \subseteq E_2$,then $P(E_2) \leq P(E_1)$

If $P(B) = \frac{3}{4}$,$P(A \cap B \cap \bar{C}) = \frac{1}{3}$ and $P(\bar{A} \cap B \cap \bar{C}) = \frac{1}{3}$,then $P(B \cap C)$ is

Four persons $A$, $B$, $C$ and $D$ throw an unbiased die, turn by turn, in succession till one gets an even number and wins the game. What is the probability that $A$ wins if $A$ begins?

Vedclass Products

For Students

Vedclass Test Series

Mock tests in real JEE/NEET style with performance analysis. 5-day free trial.

Start Free Trial
For Teachers

Exam Paper Generator

Generate Set A/B/C/D exam papers from 7.5L+ questions in 2 minutes. 3 chapters free.

Try Free
For Institutes

Online Exam Module

Live online exams with unlimited students, 360° analytics & white-label branding.

See Demo